OpenStudy (anonymous):

I need help with this logarithmic function F(x)=log1/3 (3-x). How do I graph this?

OpenStudy (jdoe0001):

\(\bf log\left(\cfrac{1}{3}(3-x)\right) \qquad ?\)

OpenStudy (anonymous):

1/3 is the base.

OpenStudy (anonymous):

\[f(x)=\log_{1/3} (3-x)\]

OpenStudy (jdoe0001):

or \(\bf \large log_{\frac{1}{3}}(3-x)\quad ?\)

OpenStudy (jdoe0001):

ok

OpenStudy (jdoe0001):

ahemm... well, sorry for the delay, I'd say use the "change of base rule", that is \(\bf \textit{log change of base rule}\color{blue}{ log_ab = \cfrac{log_cb}{log_ca}}\\ \large log_{\frac{1}{3}}(3-x) \implies \cfrac{log_{10}(3-x)}{log_{10}\left(\frac{1}{3}\right)}\)

OpenStudy (anonymous):

Will do, thanks!

OpenStudy (jdoe0001):

yw